OpenStudy (anonymous):

Consider the curve y=e^xcosx A) find the x-coordinate of M, the maximum value of y between 0

OpenStudy (turingtest):

what have you tried so far?

OpenStudy (anonymous):

i have tried to solve for x, but the issue is that there is no way to do so since i don't have a number for why

OpenStudy (turingtest):

did you take the derivative first?

OpenStudy (anonymous):

no... i'll try that now

OpenStudy (turingtest):

if you have a function \(y=f(x)\), relative maximums and minimums occur when \(y'=f'(x)=0\)

OpenStudy (anonymous):

true... ill try that out now

OpenStudy (anonymous):

@TuringTest I found the derivative to the original equation,\[y \prime= e^x(\cos(x)-\sin(x))\], but i am not sure how i am supposed to find x once i equal it to zero